The following equation summarizes the trend portion of quarterly sales of condominiums over a long cycle. Sales also exhibit seasonal variations.

Ft = 62 − 5.2t + 4.2t 2
  

where
Ft = Unit sales
t = 0 at the first quarter of last year

Quarter Relative
1 1.15
2 1.00
3 .65
4 1.20


Using the information given, prepare a forecast of sales for each quarter of next year (not this year), and the first quarter of the year following that. (Round intermediate calculations and final answers to 2 decimal places.)
  

Quarter Forecast
1
2
3
4
1
Answer #1

Answer:

Quarter relative, r Forecast using, Ft= 62-5.2t + 4.2t^2
1 1.15 61.00
2 1 68.40
3 0.65 84.20
4 1.2 108.40

Now multiply Ft and relative factor (r)

Quarter relative, r Forecast using, Ft= 62-5.2t + 4.2t^2 Forecast with relative index= Ft*r
1 1.15 61.00 70.15
2 1 68.40 68.4
3 0.65 84.20 54.73
4 1.2 108.40 130.08